OpenStudy (anonymous):

The points scored in the highest scoring game of the season was 3/11 of the total points, T, scored during the entire season. The points scored in the second highest scoring game was 3/11 of the remainder. If these points differed by 9, what was T, the total points scored in the season?

OpenStudy (anonymous):

3/11 of the remainder?

OpenStudy (anonymous):

as in, matches xcept the highest scoring match?

OpenStudy (anonymous):

3/11 of remainder score

OpenStudy (anonymous):

okay so total score = T the highest scoring match's score = 3T /11

OpenStudy (anonymous):

and the 2nd highest scoring game 3/11 of the remainder,which is 3/11 of T - 3T/11 right?

OpenStudy (anonymous):

thats 3/11 * 8T/11

OpenStudy (anonymous):

thats 24 T /121

OpenStudy (anonymous):

3T/11 - 24T /121 = 9

OpenStudy (anonymous):

so find T do u hv options for answer?

OpenStudy (anonymous):

so t is 121 right ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

what are the answers given then?

OpenStudy (anonymous):

@BSwan where did i go wrong here

OpenStudy (anonymous):

i think i didnt get the remainder part

OpenStudy (anonymous):

ok so you would have equation:- \(\Huge \frac{3T}{11}- \frac{3}{11}(\large {T}- \Huge \frac{3T}{11})=9\)

OpenStudy (anonymous):

so @aajugdar did u reach this equation ?

OpenStudy (anonymous):

i did ,look above

OpenStudy (anonymous):

she says its incorrect pffft

OpenStudy (anonymous):

ur answer is correct :) @aajugdar

OpenStudy (anonymous):

well lets c what she has to say

OpenStudy (anonymous):

thanks @aajugdar

OpenStudy (anonymous):

its correect

OpenStudy (anonymous):

-.- yaw